OpenStudy (anonymous):

what is x in the square root of the square root of 4= 2^x

OpenStudy (anonymous):

 is this the ques??

hartnn (hartnn):

or this ? \(\huge \sqrt{\sqrt{4}}=2^x\) ??

OpenStudy (anonymous):

hartnn yes thats it

hartnn (hartnn):

so write 4 as \(2^2\)

OpenStudy (anonymous):

sqrt 4 = 2 so use this

hartnn (hartnn):

after writing √4=2 u have \(\sqrt 2\) which u can write as \(2^{1/2}\) which equals 2^x so now can u tell me whats x ?

hartnn (hartnn):

as 2^(1/2)=2^x x=1/2